# Fast sampling and model selection for Bayesian mixture models

### v1

@@ Abstract
We describe two Monte Carlo algorithms for sampling from the integrated posterior distributions of a range of Bayesian mixture models. Both algorithms allow us to directly sample not only the assignment of observations to components but also the number of components, thereby fitting the model and performing model selection over the number of components in a single computation. The first algorithm is a traditional collapsed Gibbs sampler, albeit with an unusual move-set; the second builds on the first, adding rejection-free sampling from the prior over component assignments, to create an algorithm that has excellent mixing time in typical applications and outperforms current state-of-the-art methods, in some cases by a wide margin. We demonstrate our methods with a selection of applications to latent class analysis.
I Introduction
A mixture model describes data or observations that are divided into a number of groups or components, with the distribution of the data being different in each component MP00 ; MLR19 . Gaussian mixture models, for instance, describe data drawn from a combination of Gaussians and fitting to such model allows one to classify data points optimally into Gaussian clusters Reynolds09 ; NF07 . Models can also be defined using mixtures of beta, gamma, Dirichlet, Poisson, or binomial distributions, or exponential families such as the generalized linear model, among many possibilities MP00 ; MLR19 . In a somewhat different context, latent class models describe data with categorical variables, such as responses to tests or surveys, and fitting allows one to identify commonly occurring patterns of responses Goodman74 ; McCutcheon87 ; BR93 ; LLSL18 . Latent profile models serve a similar purpose for real-valued data CL10 .
Fits to mixture models can be performed using standard maximum-likelihood or maximum a posteriori (MAP) methods, or more often using an expectation-maximization (EM) algorithm that returns MAP estimates of the parameters of the probability distributions characterizing each component and a complete posterior distribution over the component assignments DLR77 ; McCutcheon87 ; MP00 ; LLSL18 ; MLR19 . These methods however have some shortcomings. For instance, they return only point estimates of the parameters and can produce overfitting of the data or underdetermination of the parameter values when the number of parameters is large. But perhaps the primary issue is that they provide no direct way of estimating the number of components, a crucial part of the inference process Everitt88 ; LD97 ; RD06 ; NAM07 ; Hoijtink10 ; MR14 .
@@ I Introduction
An alternative approach is to integrate the probability parameters out of the model to yield a Bayesian marginal posterior distribution over the component assignments alone. This approach is not widely used—for example, the otherwise comprehensive recent review of MR14 fails to mention it even in passing—but it is potentially promising for practical, robust mixture model analysis. Current implementations, however, again have shortcomings. The integration is most commonly approximated numerically by Gibbs sampling WM14 ; LLSL18 , which can be computationally efficient but again does not provide a direct route for estimating the number of components: the number of components cannot itself be sampled because the size of the parameter space is not constant when the number of components is varying. Reversible jump Monte Carlo Green95 ; RG97 can get around this issue, but at some cost in code complexity and computational effort. More commonly, one just performs repeated analyses with different (fixed) numbers of components using Gibbs sampling, then selects among them using, for example, likelihood ratios or the Bayesian information criterion Everitt88 ; LD97 ; RD06 ; NAM07 ; Hoijtink10 . This is a costly procedure, however, since it requires repeated Monte Carlo runs, with most of the data being discarded in the final analysis.
An alternative is to marginalize over the probability parameters analytically, which can be done in closed form for many of the most common mixture models, using suitable conjugate priors NF07 ; BCG10 ; WWM16 ; NR16 ; Porteous08 . Then one can sample directly from the integrated posterior using one of several proposed sampling algorithms, such as the allocation sampler of Nobile and Fearnside NF07 or the modified Gibbs sampler of Porteous et al. Porteous08 . It is a version of this approach that we employ in this paper.
@@
Standard mixture models, as commonly formulated, also suffer from a technical, but important, difficulty: the existence of empty components. In many models, particularly those using a Dirichlet-categorical prior on the component assignments, the number of observations in a component can be zero. Arguably this is acceptable for a model with a fixed number of components, but when the number of components is a free random variable it causes ambiguity, because a given division of observations into components can be represented in more than one way in the model. For instance, we could divide observations into two components, or we could divide them into three components, one of which is empty. This in turn creates difficulties when estimating the number of components—do we have two components or three? One proposal for avoiding this problem is to use a Dirichlet-process prior such as the Chinese restaurant process, which automatically disallows empty components, but the Chinese restaurant process is somewhat specialized and we prefer a more flexible choice.
@@
Our contributions in this paper are two-fold. First, we advocate in favor of a variant prior on the component assignments, closely similar to the traditional uninformative Dirichlet-categorical prior, but differing from it in that it forbids the existence of empty components. Second, and more importantly, we describe two simple but effective Monte Carlo algorithms for sampling from the resulting integrated posterior. The first is a relatively standard collapsed Gibbs sampler that includes additional Monte Carlo moves for sampling the number of components as well as the component assignment. The second builds on this approach and exploits the particular structure of the prior to create a more efficient algorithm that significantly outperforms other state-of-the-art methods. Our approach can be applied to any mixture model in principle, provided a closed-form expression is available for the integrated posterior. This includes many of the most widely used models, such as Gaussian, beta, and gamma mixture models, latent Dirichlet analysis, latent class analysis, latent profile analysis, and more exotic models such as the stochastic block model and its variants HLL83 ; ABFX08 ; KN11a .
@@
Crucially, the algorithms we describe do not fix the number of components, but allow it to vary freely, being sampled along with the component assignments themselves. This allows one trivially to determine the posterior distribution over the number of components and hence to infer the number of components as an integral part of the calculation. No additional model selection procedure, for instance using the BIC, is necessary, and only a single run of the algorithm is needed to determine both the number of components and the assignment of observations to components. Other algorithms have been proposed with similar aims, such as reversible jump Monte Carlo Green95 ; RG97 and the allocation sampler NF07 , but these methods are more complex to implement and their performance is substantially inferior to the methods reported here.
@@
As an example of our methods, we apply them to the problem of latent class analysis on a range of example data sets, both synthetic and empirical. We provide code implementing these calculations in the C and Python programming languages. Benchmark tests show that the code is fast, performing millions of Monte Carlo steps per second on standard hardware (circa 2024), which is efficient enough that calculations on typical data sets can be completed in a few seconds, substantially faster than competing Monte Carlo methods, which can take minutes to hours BCG10 ; Walsh06 .
II Mixture models
@@
Consider a data set consisting of $N$ observations $x_{i}$ with $i=1\ldots N$, where an observation can be a single number or categorical variable, or a (possibly heterogenous) collection of such variables. We assume the data are drawn from a mixture model defined as follows.
@@
First, each observation is assigned to one of a number of groups or classes, generically called components. This is sometimes done using a dynamical urn process such as the Chinese restaurant process, but more often one uses a Dirichlet-categorical distribution. First, we pick the number $k$ of components, which can take any value between 1 (all observations in a single component) and $N$ (every observation the sole member of its own component). For instance, one could use a uniform (uninformative) prior on $k$, which means
@@
Other choices are also possible: if one had a particular belief about the expected value of $k$, for instance, then the maximum-entropy (least informative) prior would be a geometric distribution $P(k)\propto a^{k}$ for some $a<1$. In our example applications we use a uniform prior on $k$, but we derive the theory for general $P(k)$.
Given $k$, each observation $i$ is assigned to a component $z_{i}=1\ldots k$ drawn from a $k$-dimensional symmetric Dirichlet-categorical distribution with some concentration parameter $\alpha$:
@@
Most often $\alpha=1$, which corresponds to a flat prior over the Dirichlet parameters, in which case
$P(z|k)={(k-1)!\over(N+k-1)!}\prod_{r=1}^{k}n_{r}!$ (3)
where $n_{r}$ is the number of observations in component $r$ and $z$ here is shorthand for the entire set of assignments $z_{i}$. An alternative way of obtaining the same result is just to note that when $\alpha=1$ all choices of the component sizes $n_{r}$ have equal probability and that the particular assignment of observations to components is uniform over the set of possible assignments conditioned on the values of the $n_{r}$. There are (${N+k-1\choose k-1}$ choices of the $n_{r}$ (following the standard stars-and-bars argument) and a multinomial $N!/\prod_{r}n_{r}!$ ways of assigning the observations, so the total probability of each assignment is
$P(z|k)={N+k-1\choose k-1}^{\!-1}{\prod_{r}n_{r}!\over N!}={(k-1)!\over(N+k-1)!}\prod_{r=1}^{k}n_{r}!$ (4)
which recovers (3).
@@
The last part of the model is the assignment of the actual values of the observations $x_{i}$. Before getting to that, however, we point out a shortcoming of the distribution in Eq. (4), that it can result in empty components. There is nothing to stop $n_{r}$ from being zero, and indeed correct normalization requires that it must be zero some portion of the time, but it is unclear what this means. What does it mean to have three components, one of which is empty? How is that different from two non-empty components? Unless the value of $k$ is known a priori, which we here assume it is not, there is no practical difference between these situations, and yet they appear as distinct component assignments in the model. As discussed in the introduction, this ambiguity causes problems when estimating the number of components $k$.
These issues can be disposed of, however, by an easy modification. We simply forbid empty components. That is, we consider only choices of the component sizes $n_{r}$ such that none of them is zero. There are (${N-1\choose k-1}$ such choices and, assuming once again that they each occur with equal frequency, the probability $P(z|k)$ becomes
@@
This is the choice we use in this paper. Our methods can be applied also to the (more conventional) choice in Eq. (4), but we consider (5) a more satisfactory version of the model.
@@
II.1 Fitting the model
At this point, one commonly proceeds in one of two ways. One is to fit the model to the data using an EM algorithm, but for the reasons outlined in the introduction we prefer a Bayesian approach. The most common version uses a Gibbs sampler which, for given $x$ and $k$, samples alternately from the marginal distributions of the parameters $\theta_{r}$ and the component assignments $z_{i}$ Gelfand00 ; WM14 ; WWM16 . Discarding the $\theta_{r}$ then gives us a true sample from the marginal posterior $P(z|x,k)$, which can be used to estimate component assignments. This method does not however give us a direct way of estimating $k$—typically $k$ is instead determined by repeating the entire calculation for various values of $k$ and then selecting between them using, for instance, the Bayesian information criterion RD06 ; MR14 . This adds extra work and complexity to the calculation. Extensions of the Gibbs sampling approach have been proposed that allow for direct estimation of $k$ Green95 ; RG97 , but here we take a different approach. We integrate the parameters $\theta$ out of the likelihood exactly and then sample directly from the resulting marginal distribution. We write
@@ II.1 Fitting the model
$P(x|k,z)=\int P(x|k,z,\theta)P(\theta|k)\>\mathrm{d}\theta.$ (9)
With the appropriate conjugate prior for $P(\theta|k)$, the integral can be completed in closed form for many of the most common mixture models. This approach allows us, in principle at least, to make estimates of $k$ as well as the component assignments $z$, because $k$ is now a free parameter of the distribution and we are at liberty to estimate it just as we would any other parameter, for instance by sampling its value. In particular, the dimension of the parameter space does not vary with $k$, as it does in the model before integration, so no variable-dimension sampling method, such as reversible jump Monte Carlo, is required.
In practice, however, implementing this approach is not as simple as it sounds. One might imagine that one could sample directly from (8) using Metropolis-Hastings Monte Carlo but, while this can be done, it typically has low numerical efficiency. A better approach is to use a so-called collapsed Gibbs sampler, although most such samplers do not sample values of $k$ and hence a separate model selection step is still needed. An exception is the allocation sampler of NF07 for Gaussian mixture models and its extension to other models WWM16 , which does sample $k$, but relies on a Metropolis-Hastings style rejection algorithm to do so, which again has low efficiency.
In this paper we describe a method for sampling directly and efficiently from the marginal posterior, Eq. (8). Our approach employs a combination of techniques to create an algorithm that is both simple and gives results of quality comparable to or better than previous approaches in significantly shorter running times.
III Monte Carlo algorithms
In this section we describe two Monte Carlo algorithms for sampling from the joint marginal posterior on $k$ and $z$, Eq. (8). The first algorithm is a relatively traditional collapsed Gibbs sampler; the second introduces an additional idea that improves efficiency.
III.1 Algorithm 1
The first algorithm we describe is a collapsed Gibbs sampler that samples both the component assignments $z$ and the number of components $k$ at the same time, allowing one to perform model fitting and model selection in a single step. The algorithm is similar in spirit to samplers for Dirichlet-process models such as those of Das Das14 and Khoufache et al. Khoufache23 . It is both simple and efficient, although in Section III.2 we describe another algorithm that is more efficient and, arguably, even simpler.
Given any initial choice of $k$ and assignment $z$ of the $N$ observations to components, our algorithm repeatedly performs the following steps. {enumerate*}
Choose an observation $i$ uniformly at random. Let $r$ be the component to which $i$ currently belongs.

### v2

@@ Abstract
We study Bayesian estimation of mixture models and argue in favor of fitting the marginal posterior distribution over component assignments directly, rather than Gibbs sampling from the joint posterior on components and parameters as is commonly done. Some previous authors have found the former approach to have slow mixing, but we show that, implemented correctly, it can achieve excellent performance. In particular, we describe a new Monte Carlo algorithm for sampling from the marginal posterior of a general integrable mixture that makes use of rejection-free sampling from the prior over component assignments to achieve excellent mixing times in typical applications, outperforming standard Gibbs sampling, in some cases by a wide margin. We demonstrate the approach with a selection of applications to Gaussian, Poisson, and categorical models.
1 Introduction
Mixture models provide a powerful framework for model-based clustering, in which data are partitioned into clusters or components, each with their own characteristic distribution, and model fits return an estimate of both the partition and the individual distributions [57, 34, 32, 22]. The fit is often performed using an expectation-maximization (EM) algorithm that returns maximum likelihood or maximum a posteriori (MAP) estimates of model parameters and a complete posterior distribution over the assignment of data to components [12, 34, 17, 36]. EM algorithms however have some shortcomings. In addition to returning only point estimates of the parameters, they can lead to overfitting of the data or underdetermination of the parameters when the number of parameters is large or the data are sparse, and they provide no direct way of estimating the number of components—commonly, one just performs repeated fits with different numbers of components, then selects among them using, for example, likelihood ratios or the Bayesian information criterion, but this a costly and cumbersome procedure, making it tempting to neglect this important step of the analysis [13, 31, 6, 49, 46, 25, 35, 40].
@@ 1 Introduction
Alternatively, we can take a Bayesian approach and impose a prior on the parameters then integrate them out of the model to yield a marginal posterior over the component assignments alone. This obviates any problems with point estimates of the parameters, but it has its own shortcomings, at least as typically implemented. The integration is most commonly approximated numerically by Gibbs sampling [3, 16, 41, 14, 22], which can be computationally efficient but again does not provide a direct route for estimating the number of components: the number cannot itself be sampled because the size of the parameter space is not constant when the number of components is varying. Reversible jump Monte Carlo and similar methods [23, 47, 50, 56] can get around this issue, although they are more computationally demanding and have been described as “notoriously hard to tune and may lead to poor mixing” [4]. More often, one again just performs repeated analyses with various (fixed) numbers of components and selects among the resulting fits, but this can also be numerically costly, since it requires repeated Monte Carlo runs, with most of the data being discarded in the final analysis.
An alternative is to marginalize analytically over the model parameters, which can be done in closed form for many of the most common mixture models, using suitable priors [44, 54, 5, 43, 62]. Then one can sample directly from the integrated posterior using one of several proposed sampling schemes, such as the allocation sampler of Nobile and Fearnside [45] or the modified Gibbs samplers of Jain and Neal [26] and Porteous et al. [48]. This approach is not widely used—for example, the otherwise comprehensive review of [35] fails to mention it even in passing—but it is potentially promising for practical, robust mixture modeling because it again removes any issues with point estimates and it allows one to sample not only the component assignments but also the number of components, and hence estimate both at the same time without the need for a separate model selection procedure. Some authors have suggested that it may suffer from slow mixing [45, 37], but we argue that, when implemented correctly, it can be highly numerically efficient. It is a version of this approach that we study in this paper.
@@
Standard mixture models, as commonly formulated, also suffer from a technical, but important, difficulty: the existence of empty components [53]. In many models the number of observations in a component can be zero, and while this is arguably acceptable for a model with a fixed number of components, when the number of components is a free random variable it generates ambiguity, because a given division of observations into components can be represented in more than one way in the model. For instance, we could divide observations into two components, or we could divide them into three components, one of which is empty. This in turn creates difficulties when estimating the number of components—do we have two components or do we have three? (We note that some authors use the word “cluster” to refer to a non-empty component, although others use the words cluster and component interchangeably. In this paper we will specifically use the phrase “non-empty component” for the sake of clarity.)
One solution to the problem of empty components is to employ a sparse model with a prior that penalizes empty components and then count only non-empty ones [58, 15]. Another is to use an infinite model with a Dirichlet-process prior and again count only non-empty components [41, 24]. Here we argue for a third approach that straddles the boundary between infinite mixtures and sparse finite ones.
@@
Our contributions are two-fold. First, we advocate in favor of a variant prior on the component assignments, closely similar to a traditional Dirichlet-categorical prior, but differing in that it forbids the existence of empty components. Second, and more importantly, we describe a Monte Carlo algorithm for sampling from the resulting integrated posterior that exploits the structure of the prior to create a more efficient algorithm that significantly outperforms other state-of-the-art methods. Our approach can in principle be applied to any mixture model for which a closed-form expression exists for the integrated posterior. Models falling in this category include many of the most widely used models, such as univariate and multivariate Gaussian mixtures with normal/gamma priors (normal distribution over the mean and gamma distribution over the inverse variance or precision); Gaussian mixtures with normal/inverse-gamma priors (inverse-gamma distribution over the variance); Dirichlet-process Gaussian models with normal/Wishart priors or normal/inverse-Wishart priors; Poisson mixtures with gamma priors; geometric mixtures with gamma priors; multivariate Bernoulli and categorical models; and binomial or multinomial mixtures with beta or Dirichlet priors.
@@
Crucially, the algorithm we describe does not fix the number of components, but allows it to vary freely, being sampled along with the component assignments themselves. This allows one trivially to determine the posterior distribution over the number of components and hence to infer the number of components as an integral part of the calculation. No additional model selection procedure, for instance using the Bayesian information criterion, is necessary, and only a single run of the algorithm is needed to determine both the number of components and the assignment of observations to components. We show that this approach is consistently able to recover the correct number of components in synthetic tests.
@@
As examples of our methods, we apply them to Gaussian, Poisson, and categorical mixture models on a range of example data sets. We provide implementations in the C and Python programming languages that can perform millions of Monte Carlo steps per second on standard hardware, sufficient that calculations on typical data sets can be completed in a few seconds, and we give timing results showing that this is substantially faster than competing Monte Carlo methods.
2 Bayesian mixture models
@@
Consider a data set of $N$ observations $x_{i}$ with $i=1\ldots N$, where an observation can be a single number or categorical variable, or a (possibly heterogeneous) collection of such variables. We assume the data are drawn from a mixture model defined as follows.
@@
First, we pick the number $k$ of components, which can take any value between 1 (all observations in a single component) and $N$ (every observation the sole member of its own component). For instance, one could use a uniform (uninformative) prior on $k$, which means
@@
Other choices are also possible: if one had a particular belief about the expected value of $k$, for instance, then the maximum-entropy (least informative) prior would be a geometric distribution $P(k)\propto a^{k}$ for some $a<1$. In our example applications we use a uniform prior on $k$, but we derive the theory for general $P(k)$. In this respect our approach offers more flexibility than, for example, models with a Dirichlet process prior, which imposes a specific choice of $P(k)$.
Given $k$, each observation $i$ is assigned to a component $z_{i}=1\ldots k$. In the first instance, let us suppose the components are drawn from a $k$-dimensional symmetric Dirichlet-categorical distribution with some concentration parameter $\alpha$:
@@
That is, we choose a set of $k$ probabilities $\pi_{r}$ with $r=1\ldots k$ from the symmetric Dirichlet distribution
$P(\pi|k,\alpha)={1\over\mathrm{B}(\alpha,\alpha,\ldots)}\prod_{r=1}^{k}\pi_{r}^{\alpha-1},$ (3)
where
$\mathrm{B}(x_{1},x_{2},\ldots)={\prod_{r}\Gamma(x_{r})\over\Gamma\bigl(\sum_{r}x_{r}\bigr)}$ (4)
is the multivariate beta function and $\pi$ is shorthand for the entire set of probabilities $\pi_{r}$. Now we choose each $z_{i}$ independently from the categorical distribution with probabilities $\pi_{r}$, so that the probability of a particular assignment of components for all $N$ observations is
$P(z|\pi)=\prod_{i=1}^{N}\pi_{z_{i}}=\prod_{r=1}^{k}\pi_{r}^{n_{r}},$ (5)
where $n_{r}$ is the number of observations in component $r$. Marginalizing over the probabilities, we then get
$P(z|k,\alpha)$ $=\int P(\pi|k,\alpha)P(z|\pi)\>\mathrm{d}\pi={1\over\mathrm{B}(\alpha,\alpha,\ldots)}\int\prod_{r=1}^{k}\pi_{r}^{n_{r}+\alpha-1}\>\mathrm{d}\pi$
$={\mathrm{B}(n_{1}+\alpha,n_{2}+\alpha,\ldots)\over\mathrm{B}(\alpha,\alpha,\ldots)}={\Gamma(k\alpha)\over\Gamma(N+k\alpha)}\prod_{r=1}^{k}{\Gamma(n_{r}+\alpha)\over\Gamma(\alpha)}.$ (6)
@@
Given these component assignments, we now generate the values of the observations $x_{i}$ themselves. Before getting to that, however, we point out a shortcoming of the distribution in Eq. (6), that it can result in empty components, as mentioned in the introduction. There is nothing to stop $n_{r}$ from being zero, and indeed correct normalization requires that it must be zero some portion of the time, but it is unclear what this means. What does it mean to have three components, one of which is empty? How is that different from two non-empty components? Unless the value of $k$ is known a priori, which we here assume it is not, there is no practical difference between these situations, and yet they appear as distinct component assignments in the model [53, 15].
These issues can be disposed of, however, by an easy modification. We simply forbid empty components, meaning we consider only component assignments where each component contains at least one observation. We can achieve this by a generative process in which we first assign one observation, chosen uniformly at random, to each of the $k$ components, then assign the remainder following a Dirichlet-categorical distribution with concentration parameter $\alpha$. There are N!/(Nk)!N!/(N-k)! ways to choose the initial $k$ assignments and probability $\prod_{r}\pi_{r}^{n_{r}-1}$ of a particular assignment of the remaining $n_{r}-1$ members of each component. Because any of the members of a component could be chosen as the first member, there are also $\prod_{r}n_{r}$ ways to generate the same complete set of assignments, so the total probability of any particular assignment is
$P(z|\pi)={(N-k)!\over N!}\prod_{r=1}^{k}n_{r}\pi_{r}^{n_{r}-1},$ (7)
and
$P(z|k,\alpha)$ $=\int P(\pi|k,\alpha)P(z|\pi)\>\mathrm{d}\pi={1\over\mathrm{B}(\alpha,\alpha,\ldots)}{(N-k)!\over N!}\prod_{r=1}^{k}n_{r}\int\prod_{r=1}^{k}\pi_{r}^{n_{r}+\alpha-2}\>\mathrm{d}\pi$
$={(N-k)!\over N!}\Biggl[\prod_{r=1}^{k}n_{r}\Biggr]{\mathrm{B}(n_{1}+\alpha-1,n_{2}+\alpha-1,\ldots)\over\mathrm{B}(\alpha,\alpha,\ldots)}$
$={(N-k)!\over N!}{\Gamma(k\alpha)\over\Gamma(N+k(\alpha-1))}\prod_{r=1}^{k}n_{r}{\Gamma(n_{r}+\alpha-1)\over\Gamma(\alpha)}$
$={1\over N!}(N-k)\,\mathrm{B}(N-k,k\alpha)\prod_{r=1}^{k}n_{r}{\Gamma(n_{r}+\alpha-1)\over\Gamma(\alpha)}.$ (8)
This is the prior we use in this paper. Our methods can be applied also to the (more conventional) form in Eq. (6), but we consider (8) a more satisfactory version of the model from both a theoretical and practical point of view.
Our approach can be applied for any value of the concentration parameter $\alpha$, but the most common choice is $\alpha=1$, which corresponds to a uniform distribution over the Dirichlet probabilities $\pi_{r}$ and hence also over the sizes $n_{r}$ of the components. For this $\alpha$, Eq. (8) simplifies to
@@
This is the choice we make for the example calculations in this paper, although our algorithms work for any value of $\alpha$.
@@
2.1 Fitting the model
At this point, one commonly proceeds in one of two ways. The first is to fit the model to the data using an EM algorithm, but for the reasons outlined in the introduction we prefer a Bayesian approach. The most common Bayesian method uses a Gibbs sampler which, for given $x$ and $k$, samples alternately from the marginal distributions of the parameters $\theta_{r}$ and the component assignments $z_{i}$ [16, 41, 22]. Discarding the $\theta_{r}$ then gives us a true sample from the marginal posterior $P(z|x,k)$, which can be used to estimate component assignments. This method does not however give us a direct way of estimating kk—typically $k$ is instead determined by selecting among models with various fixed values of $k$ using, for instance, the Bayesian information criterion [49, 35]. This adds extra work and complexity to the calculation. Extensions of the Gibbs sampling approach have been proposed that allow for direct estimation of $k$, such as reversible jump Monte Carlo [23, 50], but here we take a different approach. We integrate the parameters $\theta$ out of the likelihood exactly and then sample directly from the resulting marginal distribution. We write
@@ 2.1 Fitting the model
$P(x|k,z)=\int P(x|k,z,\theta)P(\theta|k)\>d\theta.$ (13)
With an appropriate prior $P(\theta|k)$ the integral can be completed in closed form for many of the most common mixture models. This approach allows us, in principle at least, to make estimates of $k$ as well as the component assignments $z$, because $k$ is now a free parameter of the distribution and we are at liberty to estimate it just as we would any other parameter, for instance by sampling its value. In particular, the dimension of the parameter space does not vary with $k$, as it does in the model before integration, so no variable-dimension sampling method, such as reversible jump Monte Carlo, is required.
In practice, however, implementing this approach is not trivial. One might imagine that one could sample directly from (12) using Metropolis-Hastings Monte Carlo for instance, but, while this can be done, it typically has low numerical efficiency. A better approach is to use a so-called collapsed Gibbs sampler, although most such samplers do not sample values of $k$ and hence a separate model selection step is still needed. Exceptions include the “split-merge” method of [26] and the allocation sampler of Nobile and Fearnside [45] for Gaussian mixture models and its extension to other models [62]. These methods do sample $k$, although the former is computationally demanding and the latter relies on a Metropolis-Hastings style rejection algorithm which again has low efficiency.
In this paper we describe a method for sampling directly and efficiently from the marginal posterior of Eq. (12). Our approach employs a combination of techniques to create an algorithm that is both simple to implement and gives results comparable to previous approaches in significantly shorter running times.
3 Monte Carlo algorithm
In this section we describe the proposed Monte Carlo algorithm for sampling from the joint marginal posterior on $k$ and $z$, Eq. (12), with the uniform Dirichlet-categorical prior of Eq. (9). The algorithm also extends to the more general non-uniform prior of Eq. (8) and we describe this extension in Appendix C, but here we focus on the simpler uniform case for the sake of pedagogical clarity.
The algorithm is a form of collapsed Gibbs sampler similar in spirit to samplers for Dirichlet-process models such as those of Das [10] and Khoufache et al. [27], but applied to the Dirichlet-categorical prior and with some additional innovations that improve performance. Given any initial choice of $k$ and assignment $z$ of the $N$ observations to components, the algorithm repeatedly performs the following actions.
  1. 1.
  2. 2.